# Regret Minimization in Stackelberg Games with Side Information

### v2

@@ 4.1 Stochastic Follower Types and Adversarial Contexts
We now show that the leader can obtain a complementary regret bound of $O(A_{f}\sqrt{T\log T})$ if they instead estimate the probability that the follower best-responds with action $a_{f}\in\mathcal{A}_{f}$, given a particular context $\mathbf{z}$ and leader mixed strategy $\mathbf{x}$ [^9] In general $K$ has no dependence on $A_{f}$, and vice versa. In what follows, we use $\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a_{f})}\in\{0,1\}^{K}$ to refer to the indicator vector whose $i$-th component is $\mathbbm{1}\{\sigma^{(\mathbf{z},\mathbf{x})}(\alpha^{(i)})=a_{f}\}$, i.e. the indicator that a follower of type $\alpha^{(i)}$ best-responds to context $\mathbf{z}$ and mixed strategy $\mathbf{x}$ by playing action $a_{f}$.
Corollary 4.8.
If $\;\widehat{\mathbf{p}}_{t}(\mathbf{z},\mathbf{x})=\{\widehat{p}_{t}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a_{f})})\}_{a_{f}\in\mathcal{A}_{f}}$, $\widehat{p}_{t+1}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a)})=\frac{1}{t}\sum_{\tau=1}^{t}\mathbbm{1}\{b_{f_{\tau}}(\mathbf{z},\mathbf{x})=a\}$, and $\widehat{p}_{1}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a)})=\frac{1}{A_{f}}$ for all $a_{f}\in\mathcal{A}_{f}$, then the expected regret of Algorithm 1 is bounded as
$\mathbb{E}[R(T)]=O\left(A_{f}\sqrt{T\log(T)}\right).$
The proof of Corollary 4.8 is similar to the proof of Corollary 4.7, and so we relegate it to the Appendix. Importantly, $\widehat{p}_{t+1}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a)})$ is computable for any $(\mathbf{z},\mathbf{x})$ because we observe $f_{t}$ after round $t$.
4.2 Stochastic Contexts and Adversarial Follower Types
Definition 4.9 (Expected Contextual Stackelberg Regret, II).
Input: Set of weights $\Omega$
[…]
for $t=1,\ldots,T$ do
      Sample $\pi_{t}\sim\mathbf{p}_{t}$, $a_{l,t}\sim\pi_{t}(\mathbf{z}_{t})$.
      Receive utility $u(\mathbf{z}_{t},a_{l,t},b_{f_{t}}(\pi_{t}(\mathbf{z}_{t})))$ and observe follower type $f_{t}$.
[…]
      end for
ALGORITHM 2 Learning with stochastic contexts: full feedback
Lemma 4.10.
Proof.
For any fixed policy $\pi$,
Theorem 4.11.
If $\Omega=\{\boldsymbol{\omega}\;:\;\boldsymbol{\omega}\in\Delta^{K},T\cdot\boldsymbol{\omega}[i]\in\mathbb{N},\forall i\in[K]\}$ and $\eta=\sqrt{\frac{\log\Pi}{T}}$, then Algorithm 2 obtains expected contextual Stackelberg regret (Definition 4.9)
$\mathbb{E}[R(T)]=O\left(\sqrt{KT\log T}+K\right).$
Proof.
By Lemma 4.4,
To conclude, it suffices to bound the discretization error, as
which follows from applying the standard regret guarantee of Hedge (Lemma A.1 in the Appendix). Applying Lemma 4.10,
5 Extension to Bandit Feedback
All of our results in Section 4 crucially rely on the assumption that the leader gets to observe the follower’s type after each round. However this assumption may not always hold in real-world Stackelberg settings of interest. For example, in cyber security settings it may be hard to deduce the organization responsible for a failed cyber attack. In wildlife protection domains, a very successful poacher may never be seen by the park rangers. Instead, the leader may only be able to observe the action the follower takes at each round. Following previous work, we refer to this type of feedback as bandit feedback. What can we say about the leader’s ability to learn under bandit feedback when there is side information?
While our impossibility result of Section 3 immediately applies to this more challenging setting, our algorithms from Section 4.1 and Section 4.2 do not. This is because we can no longer compute quantities such as $\mathbbm{1}\{f_{t}=\alpha^{(i)}\}$ or $b_{f_{t}}(\mathbf{z}_{t},\mathbf{x})$ for an arbitrary mixed strategy $\mathbf{x}$ from just follower $f_{t}$’s action alone. We still assume that the follower is one of $K$ different types, although $f_{t}$ is now never revealed to the leader.
We allow ourselves two relaxations when designing learning algorithms which operate under bandit feedback. First, while the leader’s utility function may still depend on the context $\mathbf{z}_{t}$, we assume that the follower’s utility is a function of the leader’s mixed strategy $\mathbf{x}_{t}$ alone, i.e. $u_{f}(\mathbf{z},\mathbf{x},a_{f})=u_{f}(\mathbf{x},a_{f})$ for all $\mathbf{z}\in\mathcal{Z}$. This allows us to drop the dependence on $\mathbf{z}_{t}$ from both the follower’s best response and the set of approximate extreme points, i.e. $b_{f}(\mathbf{z},\mathbf{x})$ becomes $b_{f}(\mathbf{x})$ and $\mathcal{E}_{\mathbf{z}}$ becomes $\mathcal{E}$. Furthermore, our definitions of contextual follower best-response region (Definition 4.1) and contextual best-response region (Definition 4.2) collapse to their non-contextual counterparts. Depending on the application domain, the assumption that only the leader’s utility depends on the side information may be reasonable. For instance, while an institution would prefer that a server with less traffic is hacked compared to one with more, a hacker might only care about the information hosted on the server (which may not be related to network traffic patterns).
Second, we design algorithms with regret guarantees which only hold against a non-adaptive adversary. However, we hypothesize that our results in this section could be extended to hold against an adaptive adversary by using more clever exploration strategies.
Despite these relaxations, the problem of learning under bandit feedback still remains challenging because of the exponentially-large size of $\mathcal{E}$. While a natural first step is to estimate $p(b_{f_{t}}(\mathbf{x})=a_{f})$ (i.e. the probability that follower at round $t$ best-responds with action $a_{f}$ when the leader plays mixed strategy $\mathbf{x}$) for all $\mathbf{x}\in\mathcal{E}$ and $a_{f}\in\mathcal{A}_{f}$, doing so naively would take exponentially-many rounds, due to the size of $\mathcal{E}$.
Building off of results in the non-contextual setting of Balcan et al. [6], we leverage the fact that the leader’s utility for different mixed strategies is not independent. Instead, they are linearly related through the frequency of follower types which take a particular action, given a particular leader mixed strategy. Therefore, it suffices to estimate this linear function (which can be done using as few as $K$ samples) to get an unbiased estimate of $p(b_{f_{t}}(\mathbf{x})=a_{f})$ for any $\mathbf{x}\in\mathcal{E}$ and $a_{f}\in\mathcal{A}_{f}$.
Borrowing from the literature on linear bandits, we use a barycentric spanner [5] to estimate $\{\{p(b_{f_{t}}(\mathbf{x})=a_{f})\}_{a_{f}\in\mathcal{A}_{f}}\}_{\mathbf{x}\in\mathcal{E}}$ in both partial adversarial settings we consider. A barycentric spanner for compact vector space $\mathcal{W}$ is a special basis such that any vector in $\mathcal{W}$ may be expressed as a linear combination of elements in the basis, with each linear coefficient being in the range $[-1,1]$.
In Section 5.1, we use the property that estimators constructed using barycentric spanners have low variance to show that an explore-then-exploit algorithm achieves $O(K^{2/3}A_{f}^{2/3}T^{2/3}\log^{1/3}T)$ expected contextual Stackelberg regret in the setting with stochastic follower types and adversarial contexts. In Section 5.2, we use the property that estimators constructed using barycentric spanners are bounded to design a reduction to our algorithm in Section 4.2 which achieves $O(KA_{f}^{1/3}T^{2/3}\log^{1/3}T)$ expected contextual Stackelberg regret whenever the sequence of contexts is chosen stochastically and the sequence of follower types is chosen by an adversary. [^12] While it may be possible to obtain $O(\sqrt{T})$ regret rates using standard estimation techniques, this would come at the cost of a linear dependence on $|\mathcal{E}|$ (and therefore an exponential dependence on $K$ and $A_{f}$) in regret.
5.1 Stochastic Follower Types and Adversarial Contexts
for $i=1,\ldots,K$ do
      for $\tau=1,\ldots,N$ do
           end for
           end for
           for $t=K\cdot N+1,\ldots,T$ do
                end for
Theorem 5.1.
$\mathbb{E}[R(T)]=O\left(K^{2/3}A_{f}^{2/3}T^{2/3}\log^{1/3}(T)\right).$
Proof Sketch.
The key step in our analysis is to show that for any best-response function $\sigma\in\Sigma$ and follower action $a_{f}\in\mathcal{A}_{f}$, $\mathrm{Var}(\widehat{p}(\mathbbm{1}_{(\sigma=a_{f})}))\leq\frac{K}{N}$ (Lemma B.1). Using this fact, we can bound the cumulative total variation distance between $\widehat{p}(\mathbbm{1}_{(\sigma^{(\mathbf{x}_{t})}=a_{f}})$ and $p(\mathbbm{1}_{(\sigma^{(\mathbf{x}_{t})}=a_{f}}))$ for any sequence of mixed strategies and follower actions in the “exploit” phase (Lemma B.3). The rest of the analysis follows similarly to the proof of Corollary 4.7. ∎
5.2 Stochastic Contexts and Adversarial Follower Types
[…]
for $\tau=1,\ldots,Z$ do
      for $t\in B_{\tau}$ do
           end for
[…]
           end for
ALGORITHM 4 Learning with stochastic contexts: bandit feedback
We now turn our attention to learning under bandit feedback when the sequence of contexts is chosen stochastically and the choice of follower type is adversarial. While we still use barycentric spanners to estimate $\{\{\widehat{p}(\mathbbm{1}_{(\sigma^{(\mathbf{x})}=a_{f})})\}_{a_{f}\in\mathcal{A}_{f}}\}_{\mathbf{x}\in\mathcal{E}}$, we can no longer do all of our exploration “up front” like in Section 5.1 because the follower types are now adversarially chosen. Instead, we follow the technique of Balcan et al. [6] and split the time horizon into $Z$ consecutive, evenly-sized blocks. In block $B_{\tau}$, we pick a random time-step to estimate $p_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$, i.e. the probability that a follower in block $B_{\tau}$ best-responds to mixed strategy $\mathbf{x}^{(\mathbf{b})}$ by playing action $a^{(\mathbf{b})}$, for every element in our barycentric spanner $\mathcal{B}$. If whenever the leader plays $\mathbf{x}^{(\mathbf{b})}$ the follower best-responds with action $a^{(\mathbf{b})}$, we set $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})=1$. Otherwise we set $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})=0$. Since the time-step in which we play $\mathbf{x}^{(\mathbf{b})}$ is chosen uniformly from all time-steps in $B_{\tau}$, $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$ is an unbiased estimate of $p_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$. While $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$ no longer has low variance since it must be recomputed separately for every block $B_{\tau}$, it is still bounded. Therefore, we can use our estimates $\{\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})\}_{\mathbf{b}\in\mathcal{B}}$, along with the corresponding linear coefficients from the barycentric spanner, to get a bounded (and unbiased) estimate for every $p(\mathbbm{1}_{(\sigma^{(\mathbf{x})}=a_{f})})$.
Theorem 5.2.
$\mathbb{E}[R(T)]\leq O\left(KA_{f}^{1/3}T^{2/3}\log^{1/3}(T)\right).$
Proof Sketch.
6 Conclusion and Discussion
We initiate the study of Stackelberg games with side information, which, despite the presence of side information in many Stackelberg game settings of interest, has so far not received attention from the community. We focus on the online setting in which the leader faces a sequence of contexts and follower types, each of which may be chosen either stochastically or by an adversary. We show that when both sequences are chosen adversarially, no-regret learning is not possible via a reduction to the online linear thresholding problem. In contrast, when either sequence is chosen stochastically, we are able to give algorithms with $O(\sqrt{T})$ regret guarantees. We also explore an extension to bandit feedback, in which we obtain $O(T^{2/3})$ regret in both settings by making use of barycentric spanners.
There are several exciting avenues for future research.
  1. 1.
    Intermediate forms of adversary. The two relaxations of the fully adversarial setting that we consider (stochastic contexts and stochastic follower types), while natural, essentially rule out the leader learning anything about the follower’s type from the context. Although we prove that learning is impossible in the fully adversarial setting, our lower bound does not rule out the existence of no-regret algorithms when, e.g. the mapping from contexts to follower types has finite Littlestone dimension. It would be interesting to explore this direction further to better pin down when no-regret learning is possible in our setting.
  2. 2.
    $O(T^{1/2})$ regret under bandit feedback. Recently Bernasconi et al. [7] obtain $O(T^{1/2})$ regret when learning in non-contextual Stackelberg games under bandit feedback against an adversarially-chosen sequence of follower types via a reduction to adversarial linear bandits [1]. A natural approach is therefore to try to apply similar ideas to our contextual setting in hopes of getting improved regret rates. However, our attempt to do so failed because applying similar steps to Bernasconi et al. results in a reduction to a generalization of the (adversarial) contextual bandit problem for which we are not aware of any regret minimizing algorithm. Nevertheless, we view exploring whether $O(T^{1/2})$ contextual Stackelberg regret is possible under bandit feedback as a natural and exciting future direction.
  3. 3.
    Computational efficiency. Our motivation for studying this problem comes from the observation that many real-world settings which are typically modeled as Stackelberg games often have some form of side information which is available to the players. It would be interesting to see if there is any empirical structure in the side information present in typical real-world Stackelberg game settings which may allow us to circumvent the existing NP-Hardness results and design algorithms with a more efficient per-round runtime.
  4. 4.
    Fully offline learning. It would be interesting to explore other variants of the problem, e.g. the fully offline setting. While our results in Section 4.1 and Section 4.2 apply to the fully offline version of our setting, it is most likely possible to obtain improved performance by considering this setting explicitly. Furthermore, it may be possible that the fully offline version of our problem is computationally tractable under certain assumptions on the data-generating process.
Appendix A Appendix for Section 4: Limiting the Power of the Adversary
See 4.4
Proof.
Observe that for any $\mathbf{z}$,
[…]
A.1 Section 4.1: Stochastic Follower Types and Adversarial Contexts
@@ Appendix A Appendix for Section 4: Limiting the Power of the Adversary
A.2 Section 4.2: Stochastic Contexts and Adversarial Follower Types
@@
Appendix B Appendix for Section 5: Extension to Bandit Feedback
B.1 Section 5.1: Stochastic Follower Types and Adversarial Contexts
Lemma B.1.
@@ Appendix B Appendix for Section 5: Extension to Bandit Feedback
By Lemma B.1 and a Hoeffding bound, we have that
@@
where the last line follows from Lemma B.3. The desired result follows by the setting of $N$. ∎
B.2 Section 5.2: Stochastic Contexts and Adversarial Follower Types
Definition B.4.
@@
where the first line uses Lemma B.5 and the fact that $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}\sim\mathcal{P}$ are i.i.d., and $R_{N,\kappa}$ is the regret of Hedge after $N$ time-steps when losses are bounded in $[-\kappa,\kappa]$. Rearranging terms and using the fact that the expected regret of Hedge after $N$ time-steps is at most $\sqrt{N\kappa\log|\Pi|}$ gets us the desired result.∎
See 5.2
@@
where the first line comes from multiplying and dividing by $|B_{\tau}|$, the second line comes from the definition of $p_{\tau}$, the third from the definition of $u_{\tau}$, the fourth follows from linearity of expectation and the fact that $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ are i.i.d., the fifth follows from applying Lemma B.6, and the sixth line follows from the definition of $B$ and the fact that $|\Pi|\leq N^{K}$. Setting $N$ gets us the final result. ∎

### v3

@@ Appendix B Appendix for Section 4: Limiting the power of the adversary
B.2 Section 4.2: Stochastic contexts and adversarial follower types
@@
See 4.10
Proof.
For any fixed policy $\pi$,
See 4.11
Proof.
By Lemma 4.4,
To conclude, it suffices to bound the discretization error, as
which follows from applying the standard regret guarantee of Hedge (Lemma B.1 in the Appendix). Applying Lemma 4.10,
Appendix C Appendix for Section 5: Extension to bandit feedback
C.1 Stochastic follower types and adversarial contexts
Theorem C.1.
$\mathbb{E}[R(T)]=O\left(K^{2/3}A_{f}^{2/3}T^{2/3}\log^{1/3}(T)\right).$
Proof Sketch.
The key step in our analysis is to show that for any best-response function $\sigma\in\Sigma$ and follower action $a_{f}\in\mathcal{A}_{f}$, $\mathrm{Var}(\widehat{p}(\mathbbm{1}_{(\sigma=a_{f})}))\leq\frac{K}{N}$ (Lemma C.3). Using this fact, we can bound the cumulative total variation distance between $\widehat{p}(\mathbbm{1}_{(\sigma^{(\mathbf{x}_{t})}=a_{f}})$ and $p(\mathbbm{1}_{(\sigma^{(\mathbf{x}_{t})}=a_{f}}))$ for any sequence of mixed strategies and follower actions in the “exploit” phase (Lemma C.5). The rest of the analysis follows similarly to the proof of Corollary 4.7. ∎
C.2 Stochastic contexts and adversarial follower types
We now turn our attention to learning under bandit feedback when the sequence of contexts is chosen stochastically and the choice of follower type is adversarial. While we still use barycentric spanners to estimate $\{\{\widehat{p}(\mathbbm{1}_{(\sigma^{(\mathbf{x})}=a_{f})})\}_{a_{f}\in\mathcal{A}_{f}}\}_{\mathbf{x}\in\mathcal{E}}$, we can no longer do all of our exploration “up front” like in Section C.1 because the follower types are now adversarially chosen. Instead, we follow the technique of Balcan et al. [5] and split the time horizon into $Z$ consecutive, evenly-sized blocks. In block $B_{\tau}$, we pick a random time-step to estimate $p_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$, i.e. the probability that a follower in block $B_{\tau}$ best-responds to mixed strategy $\mathbf{x}^{(\mathbf{b})}$ by playing action $a^{(\mathbf{b})}$, for every element in our barycentric spanner $\mathcal{B}$. If whenever the leader plays $\mathbf{x}^{(\mathbf{b})}$ the follower best-responds with action $a^{(\mathbf{b})}$, we set $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})=1$. Otherwise we set $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})=0$. Since the time-step in which we play $\mathbf{x}^{(\mathbf{b})}$ is chosen uniformly from all time-steps in $B_{\tau}$, $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$ is an unbiased estimate of $p_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$. While $\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})$ no longer has low variance since it must be recomputed separately for every block $B_{\tau}$, it is still bounded. Therefore, we can use our estimates $\{\widehat{p}_{\tau}(\mathbbm{1}_{(\sigma^{(\mathbf{x}^{(\mathbf{b})})}=a^{(\mathbf{b})})})\}_{\mathbf{b}\in\mathcal{B}}$, along with the corresponding linear coefficients from the barycentric spanner, to get a bounded (and unbiased) estimate for every $p(\mathbbm{1}_{(\sigma^{(\mathbf{x})}=a_{f})})$.
Theorem C.2.
$\mathbb{E}[R(T)]\leq O\left(KA_{f}^{1/3}T^{2/3}\log^{1/3}(T)\right).$
Proof Sketch.
C.3 Proofs for Section C.1: Stochastic follower types and adversarial contexts
for $i=1,\ldots,K$ do
      for $\tau=1,\ldots,N$ do
           end for
           end for
           for $t=K\cdot N+1,\ldots,T$ do
                end for
Lemma C.3.
@@ C.2 Stochastic contexts and adversarial follower types
By Lemma C.3 and a Hoeffding bound, we have that
@@
where the last line follows from Lemma C.5. The desired result follows by the setting of $N$. ∎
C.4 Proofs for Section C.2: Stochastic contexts and adversarial follower types
[…]
for $\tau=1,\ldots,Z$ do
      for $t\in B_{\tau}$ do
           end for
[…]
           end for
Algorithm 4 Learning with stochastic contexts: bandit feedback
Definition C.6.
@@
where the first line uses Lemma C.7 and the fact that $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}\sim\mathcal{P}$ are i.i.d., and $R_{N,\kappa}$ is the regret of Hedge after $N$ time-steps when losses are bounded in $[-\kappa,\kappa]$. Rearranging terms and using the fact that the expected regret of Hedge after $N$ time-steps is at most $\sqrt{N\kappa\log|\Pi|}$ gets us the desired result.∎
See C.2
@@
where the first line comes from multiplying and dividing by $|B_{\tau}|$, the second line comes from the definition of $p_{\tau}$, the third from the definition of $u_{\tau}$, the fourth follows from linearity of expectation and the fact that $\mathbf{z}_{1},\ldots,\mathbf{z}_{T}$ are i.i.d., the fifth follows from applying Lemma C.8, and the sixth line follows from the definition of $B$ and the fact that $|\Pi|\leq N^{K}$. Setting $N$ gets us the final result. ∎